OpenStudy (henrietepurina):

Choose the point-slope form of the equation below that represents the line that passes through the points (-3, 2) and (2, 1).

OpenStudy (henrietepurina):

y + 3 = -5(x - 2) y - 2 = -5(x + 3) y + 3 = -one fifth(x - 2) y - 2 = -one fifth(x + 3)

OpenStudy (henrietepurina):

Help for a medal!